OpenStudy (abbycross167):

Can someone help me answer this question? I'm trying to do it myself, but I would like someone to check my answer. Please? Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. 3x - 2y= 18

OpenStudy (mathmale):

Hello, Abby. You and I worked on this problem briefly a few hours ago. I saw that several others had helped you with it. Are you posting it fresh, again?

OpenStudy (abbycross167):

no sir I don't think this is the same question... this is a different lesson. I worked out this problem by myself, using the answer that @Directrix showed me on the last question

OpenStudy (abbycross167):

Here is my work for this question.... 3x - 2y= 18 I used the intercept method 3(0)-2y=18 0-2y=18 2y=18 18/2=9 The y intercept is (0,-3) Let y =0 and find x. 3x-2(0)=18 3x-0=18 3x=18 18/3=6 The x intercept is (6,0)

OpenStudy (abbycross167):

Would one of y'all mind checking over my work? please

Directrix (directrix):

The one we finished before is this one, according to my notes: 9x + 12y= -36

OpenStudy (abbycross167):

yes sir/ma'am

OpenStudy (abbycross167):

I tried to follow the steps for that one

Directrix (directrix):

There is a problem here: y-intercept 3(0)-2y=18 0-2y=18 You lost the negative sign. - 2y=18 y = 18 divided by Negative Two y = ? @abbycross167

Directrix (directrix):

This is correct: 3x-2(0)=18 3x-0=18 3x=18 x = 18/3=6 The x intercept is (6,0)

Directrix (directrix):

@abbycross167 What is the y-intercept?

OpenStudy (abbycross167):

Oh ok I see the y intercept is -9

OpenStudy (abbycross167):

thank you for checking/fixing my work!!

Directrix (directrix):

What about the graph?

Directrix (directrix):

I suppose you drew it once you had the two points. You are welcome.

OpenStudy (abbycross167):

yes sir/ma'am